83 84 85 69.99 39.19 41.18 23.36 8055 8056 8057 8058 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 10! Required information [The following information applies to the questions displayed below.) NOTE: Throughout this lab, everyA B C D н. I 1 Date 2 10/1/23 3 10/2/23 4 10/2/23 5 10/3/23 6 10/3/23 7 10/3/23 8 10/3/23 9 10/3/23 10 10/3/23 11 10/3/23 128026 8027 8028 8029 8030 42 43 74 45 46 47 18 19 50 51 52 53 54 55 56 57 58 59 50 51 52 53 54 55 56 57 58 59 70 71 72 73 74 78081 8081 8082 8083 8084 8085 8086 8087 8088 8089 8089 8089 8090 8090 8091 8092 12.06 52.53 43.51 12.06 43.51 43.51 19.76 92.165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 18136 8137 19.76 36.4 19.76 8138 8138 34.65 69.99 34.65 614.91 426.41 497.38 59.99 59.99 354.66 231.54 8139 8140 8141 8142 284Required: 1. Which customer number is the least profitable with $15 gross margin? O 3056 O 3008 O 3013 O 3070 2. What is the5. Swap Customer for SKU. Which SKU had the highest sales for Thrustmaster? O PED-TPR O FLI-USB O WHL-TGT O PCG-HOW
Answer #1

1)

Option B - 3008

Gross margin = Sales price - Cost = 34.65 - 19.76 = 14.89 or 15

2)

Option C - 137

There are total 4 transactions with customer 3080

Total Sales Price = 117.83 + 21.99 + 59.99 + 117.83 = 317.64

Total Cost = 69.36 + 12.06 + 30.17 + 69.36 = 180.95

Total Gross margin = Total Sales price - Total cost

= 317.64 - 180.95 = 136.69 or 137

3)

Option D - 15

If we swap sales price with gross margin, then 14.89 is new sales price. Rounded off to 15

4)

Option D - 3013

There are total 8 transactions done with customer 3013

Total of such sales are 79.99 + 497.38 + 426.41 + 79.99 + 614.91 + 249.99 + 97.3 + 497.48

= 2543

5)

Option C - WHL-TGT

Total transactions for SKU WHL-TGT are 15 and each transaction is for 614.91

Total Sales = 614.91 x 15 = 9223.65 or 9,224